2,3,3,5,10,13,?,43,172,177
a)23
b)38
c)39
d)40

Answer by Edwin McCravy(20086)

Start with 2, add 1, multiply by 1, add 2, multiply by 2, etc.

    2
2+1=3
3×1=3
3+2=5
5×2=10
10+3=13
13×3=39
39+4=43
43×4=172
172+5=177

Edwin